OpenStudy (anonymous):

7y^9 x (-2y^3) explain.

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

(7y^9)(-2y^3) = - 7*2 y^(9+3) = - 14 y^12

OpenStudy (anonymous):

oh okay hold on see if i do this one right

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

-7c^7 x (-6c^8) would be -7c^7 x -6c^8 which is -2503159

OpenStudy (anonymous):

@CO_oLBoY

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

what?????? no

OpenStudy (anonymous):

well i put it into the google calculator and this is what came up lol

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

-7c^7 x -6c^8 =(-7)(-6) c^(7+8) =42 c^15

OpenStudy (anonymous):

u have to add the powers

OpenStudy (anonymous):

see i always get confused, i thought you took the parenthesis away and worked it out like that

OpenStudy (anonymous):

so 2x^3 x 8x^8 would be 16^11?

OpenStudy (anonymous):

u can took the parenthesis away and put x sign instead. They both mean the same thing its upto u which ever u wanna keep

OpenStudy (anonymous):

yup its 16x^11

OpenStudy (anonymous):

and ask different ques in separate ques

OpenStudy (anonymous):

why so you can get more medals lol

OpenStudy (anonymous):

lol no coz i hate notifications popping up again and again from same que